a rectangle measures 50 yards long and 20 yards wide. if a pole is placed every 5 yards around the perimeter how many poles are needed?
Answer by Maths68(1474)


Perimeter of a Rectangle =P=?
Length = l=50 yards
Width = w = 20 yards
Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle = 2l + 2w
P=(2*50)+(2*20)
P=100+40
P=140 yards
If a pole is placed every 5 yards around the perimeter means we have to divide 140 yards by 5
Total poles required around the rectangle at a distance of 5 yards = 140/5 = 28